suppose that the functions q and r are defined as follow
Q(x)=x^2+6
R(x)= square root x+9
find the following
(roq)(7)=
(qor)(7)=

Answer by algebrahouse.com(1659)

Q(x) = x² + 6
R(x) = √(x + 9)

(roq)(7)
= √(x² + 6 + 9) {substituted (x² + 6) in for x}
= √(x² + 15) {combined like terms}
= √[(7)² + 15] {substituted 7 for x}
= √(49 + 15) {evaluated exponent}
= √64 {added}

(qor)(7)
= [√(x + 9)]² + 6 {substituted √(x + 9) for x}
= x + 9 + 6 {evalauted exponent}
= x + 15 {combined like terms}
= 7 + 15 {substituted 7 for x}
= 22 {added}

For more help from me, visit: www.algebrahouse.com